OpenStudy (anonymous):

What is the value of 8P4?

OpenStudy (anteater):

So a permutation? 8 objects taken 4 at a time. Use: n!/(n-r)! where n = 8 and r = 4

OpenStudy (anteater):

What did you get?

OpenStudy (anteater):

So, 8!/(8-4)! = 8!/4! = ?

OpenStudy (anteater):

(8 X 7 X 6 X 5 X 4 X 3 X 2 X 1)/(4 X 3 X 2 X 1) = ...